# Squared-field cross-correlation between kinetic Sunyaev-Zel’dovich effect and 21-cm intensity mapping

### v1

@@ Abstract
Neutral hydrogen (HI) 21-cm intensity mapping is an effective method to track the distribution of baryonic matter, and extract astrophysical and cosmological information. The 21-cm intensity field has a non-vanishing cross-correlation with the kinetic Sunyaev-Zel’dovich (kSZ) effect that traces the velocity and density perturbations of free electrons. By using the linear perturbation theory, in this paper we calculate analytically, for the first time, the cross-correlation between the squared kSZ field and the projection of the squared HI intensity mapping field with the flat-sky approximation. This statistic remains non-vanishing even after the long-wavelength line-of-sight modes ($k_{\parallel}$) are removed due to foreground contamination. We further forecast for the prospects of detection with the SKA-MID 21-cm intensity mapping experiments (redshifts in range of $0.3<z<1$), and the kSZ maps measured by Atacama Cosmology Telescope (ACT) and Simons Observatory (SO). The predicted cumulative signal-to-noise ratio (SNR) is $1.92$ for SKA-ACT and $3.99$ for SKA-SO. These results show a possible on-the-edge detection on the cross-correlation signal at low redshifts, which in turn could serve as a validation step towards using it for the Epoch of Reionization studies.
@@ I Introduction
where $\tau$ is the optical depth, $\sigma_{\mathrm{T}}$ is the Thomson cross-section. $\bar{n}_{\rm e}(z)$ is the mean electron density at redshift $z$, which is equal to
@@
where $\bar{n}_{\mathrm{b},0}$ is the present day’s mean number density of baryons, and $X_{\mathrm{e}}(z)$ is the mean ionized fraction at redshift $z$, $\rho_{\rm cr}=1.879h^{2}\times 10^{-29}\,{\rm g}\,{\rm cm}^{-3}$ is the critical density of the Universe, $\Omega_{\rm b}$ is the fractional baryon density, $\mu_{\rm e}\simeq 1.14$ is the mean mass per electron Ma and Zhao (2014).
@@
In recent years, researchers have considered potential cross-correlations between the patchy reionization-era kSZ signal and other tracers of the Epoch of Reionization (EoR) Hotinli and Johnson (2022); Bégin et al. (2022); Georgiev et al. (2024); Zhou et al. (2025). Among all the large surveys, neutral hydrogen (HI) 21-cm intensity mapping (IM) survey becomes one of the most powerful large-scale structure surveys that can cover wide redshift ranges and large sky areas with efficient scanning strategies Lidz et al. (2011); Battye et al. (2013); Jiang et al. (2023). In recent years, Green Bank Telescope Chang et al. (2010); Masui et al. (2013), Parkes Telescope Pen et al. (2009); Anderson et al. (2018); Tramonte et al. (2019); Tramonte and Ma (2020), and MeerKAT telescope Cunnington et al. (2022) have measured the HI IM to a higher and higher precision, and it is almost certain that the future Square Kilometer Array will provide a large survey sample of the HI intensity maps which will constrain cosmological parameters with high precision Santos et al. (2017); Bacon et al. (2020).
@@
In this paper, we consider the cross-correlation between the squared kSZ field and the squared HI intensity field. In each case, it is necessary to square the fields before cross-correlating them. For the kSZ effect, because of the statistical isotropy, clusters with positive and negative peculiar velocities have equal weights. Therefore the direct two point correlation function $\langle\delta T_{\mathrm{kSZ}}(\vec{k}_{1})\delta T_{\mathrm{HI}}(\vec{k}_{2})\rangle$ simply vanishes given enough statistics, which can be avoided by taking the square of the kSZ field Dore et al. (2004). In practice, a Wiener filter will be applied before squaring the field to suppress the primary CMB, lensing effects, residual foreground contamination, and detector noise. In the case of the 21-cm field, without squaring, the relevant signal would involve the bispectrum $\langle\delta T_{\mathrm{kSZ}}(\vec{k}_{1})\delta T_{\mathrm{kSZ}}(\vec{k}_{2})\delta T_{\mathrm{HI}}(\vec{k}_{3})\rangle$, which is proportional to $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3})$. Since the kSZ effect is a projected signal, this statistic only receives contributions with $k^{\parallel}_{1}$ and $k^{\parallel}_{2}$ nearly equal to zero, while the Dirac Delta function enforces $k^{\parallel}_{3}=0$, as required by the translational invariance. This means that only the low $k^{\parallel}$ modes in HI field can contribute to the estimator. However, because the large-scale radial mode of HI IM is heavily contaminated by the spectrally-smooth foregrounds, the HI IM data after foreground removal procedure (either through PCA Davis et al. (1985), FASTICA Wolz et al. (2014), or GNILC Dai and Ma (2025) methods) tend to lose all low-$k^{\parallel}$ modes. Therefore, we need to square the HI IM cube so that we can fully utilize the high-$k^{\parallel}$ modes that are not removed by foreground removal procedure. In Appendix A, we prove that, by taking the square of the HI IM field before projecting it along the line-of-sight we take advantage of all high-$k^{\parallel}$ modes that contribute to the cross-correlation signal. Therefore, our goal is to model cross-correlations between the kSZ-squared field and the projection of the 21-cm-squared field. This is a 2D cross-correlation which can be performed at the map level.
@@
On large scales, the kSZ effect carries imprints of ionized baryons and follows the distribution of dark matter, which is also the case for the HI field. Therefore, we expect to see a positive correlation function between the two on large scales. On small, galactic scales neutral hydrogen resides predominantly in dense, neutral regions which remain self-shielded against photoionizing radiation Rahmati et al. (2013); Liu et al. (2022). In contrast, free electrons occupy the ionized phase of the circumgalactic medium (CGM), leading to an anti-correlation between the two tracers, which could be seen if the maps have sufficient resolution. However, this small-scale anti-correlation signal may be challenge to see because of the coarse resolution of foreseeable 21-cm measurements. Therefore, we focus our effort in predicting the signal on large scales with $\ell\lesssim 900$, and forecast its detectability.
@@
The paper is organized as follows. In Section II, we provide a detailed derivation of the analytical result consisting of the projection method II.1, general formulation of cross-correlations II.2, and the triple-spectrum evaluation II.3. In Section III we show our numerical results of the cross-correlation III.1 and the signal-to-noise ratio forecasts III.2 by adopting instrumental parameters of SKA-MID and ACT/SO kSZ measurement. The Section IV is devoted to conclusions and discussions.
@@ II Cross-correlation of the projection fields
In this section, we derive an analytical equation for the cross-correlation between kSZ square field and HI intensity mapping square field. We first go through the 2D projection procedure of a 3D field under the flat-sky approximation, and then provide a detailed calculation.
@@ II.1 Two dimension projection field
where $\chi$ is the comoving distance to the observer, $W_{X}$ is the integration kernel for $X$ field satisfying the normalization condition $\int\mathrm{d}\chi W_{X}(\chi)=1$, and $B_{X}$ is the telescope beam function in the real space which is convolved with the projected field Dore et al. (2004).
@@
Figure 1: An intuitive picture and notation explanation for the flat-sky approximation, with $\chi$ as the comoving distance and $\hat{n}_{0}$ as the line-of-sight. A harmonic decomposition under flat-sky limit can be approximated to be a 2D Fourier transform over $\vec{\theta}$.
@@
where $B_{\mathrm{HI}}$ denotes the beam factor for HI intensity mapping. In real 21-cm experiments, low-$k^{\parallel}$ modes are often lost due to the foreground removal (foreground are smooth in $k^{\parallel}$ direction), so effectively a filter must be applied to the theory to mimic this observational consequence. In the stage of numerical evaluation, we will apply a low-$k^{\parallel}$ mode cut-off $k^{\parallel}_{\mathrm{min}}=0.01~\mathrm{Mpc}^{-1}$ as the effective filter. $\bar{T}_{\mathrm{b}}(\chi)$ denotes the average brightness temperature at comoving distance $\chi$, which can be evaluated as Harper et al. (2018); Battye et al. (2013)
@@
II.2 Cross-correlating kSZ square field with projection of HI square field
We define the cross-correlation power spectrum of the kSZ square field $\delta T_{\mathrm{kSZ}^{2}}$ and the projection of HI intensity mapping square field $\delta T_{\mathrm{HI}^{2}}$ as
@@ II.2 Cross-correlating kSZ square field with projection of HI square field
The four-point correlation term $\mathcal{T}\equiv\langle q_{z}q_{z}\delta\delta\rangle$ should be proportional to a Dirac delta function by momentum conservation/translation invariance, together with a triple-spectrum $\tilde{\mathcal{T}}$ factor
@@
For small angular scales $k^{\parallel}\chi\ll\ell$, one can apply the Limber approximation, allowing us to neglect any $k^{\parallel}$ dependence in the triple-spectrum as long as the parallel mode is not enrolled in the convolution, therefore the integral over $k^{\parallel}_{1}$, $k^{\parallel}_{2}$ and $k^{\parallel}_{3}$ can be performed to obtain the Dirac delta functions of comoving distances
@@
which further simplifies Dirac delta function of perpendicular modes as $\delta_{\mathrm{D}}^{(2)}\left(\vec{\ell}+\vec{\ell}^{\prime}\right)$. Integrating over $\vec{\ell}^{\prime}$ leads to $C(\vec{\ell})$ expression as follows
@@ II.3 Triple-spectrum calculation
Here we calculate the four-point correlation term appearing in triple-spectrum $\tilde{\mathcal{T}}$, or equivalently the four-point correlator $\mathcal{T}$. An important feature of the kSZ effect Kaiser (1984); Jaffe and Kamionkowski (1998); Ma and Fry (2002) is that most contributions to the signal in integral (1) comes from modes perpendicular to the line-of-sight with $\hat{n}_{0}\cdot\vec{k}=0$ because of (near) cancellations along the line-of-sight direction. Any 3-D vector field can be decomposed into a curl-free component ($E$ field) and divergence free component ($B$ field), $\vec{q}=\vec{q}_{E}+\vec{q}_{B}$, satisfying $\nabla\times\vec{q}_{E}=0$ and $\nabla\cdot\vec{q}_{B}=0$, or in Fourier space $\vec{q}_{E}\propto\hat{k}$ and $\vec{k}\cdot\vec{q}_{B}=0$. Then from the definition given in Eq. (10), the line-of-sight projection can be evaluated as
@@
where have used the fact that the dominant contribution of kSZ effect comes from $B$ field.
@@
where $\mu_{1}=\hat{n}_{0}\cdot\hat{p}_{1}$ and $\mu_{2}=\hat{n}_{0}\cdot\hat{p}_{2}$. Further assuming Gaussianity of the perturbation fields allows us to expand the six-point correlation with Wick Theorem, to a summation of fifteen terms in total (see, e.g. Table 1). Before we move into the details of calculation, by observing the underlying symmetry of the arguments in $\mathcal{T}$, some simplifications and reduction can be carried out. First, permutation of $\vec{k}_{3}$ and $\vec{k}_{4}$ does not change the integral, therefore some of the terms give exactly the same contribution, at the integrand level. For the same reason, permutation of the vector pair $(\vec{k}_{1},\vec{p}_{1})$ and $(\vec{k_{2}},\vec{p}_{2})$ will also result in the same integral because of the exchangeability between $\vec{k}_{1}$ and $\vec{k}_{2}$. As a consequence, all independent terms are shown in Table 1 explicitly, where power spectrum of any two-point correlation is given by
@@
By applying the arguments replacement (Eq.(29)), the Dirac delta function $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{3}+\vec{k}_{4})\to\delta_{\mathrm{D}}(k_{3}^{\parallel})\delta_{\mathrm{D}}^{(2)}(\vec{\ell}^{\prime}/\chi_{3})$ indicates a zero-contribution to $C_{\ell}$ because it gives out a factor proportional to $\delta_{\mathrm{D}}^{(2)}(\vec{\ell}^{\prime})$. The cross-correlation Eq. (18) has put a constraint $\vec{\ell}+\vec{\ell}^{\prime}=0$ on $\ell$, suggesting that Term $1$ will only contribute to the monopole $\ell=0$, which is not considered. The same factor $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{3}+\vec{k}_{4})$ appears in Terms $2$ and $3$, thus there is no contribution from these two terms either. Next is Term $4$:
@@
Mathematically, the resultant zero value in the above evaluation is because Term $2$ is proportional to $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{1})$, which does not pick up a specific value of $\vec{p}_{1}$ in the integral, but rather sum over all the directions, leading to a factor proportional to zero. Similarly, Term $5$ is also zero because it is proportional to $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{2})$.
Therefore, the first non-trivial contribution is from Term 66
@@
where the minus sign in the front is resulted from two copies of velocity fields with imaginary parts, as shown in Eq. (11). Note that $k^{\parallel}_{3}$ does not stand for the $k^{\parallel}_{3}$ in Eq. (20) but the $z$-component of $\vec{k}_{3}$. Applying the replacement (29), the triple-spectrum from Term $6$ reads
@@
One can perform a decomposition of convolution momentum $\vec{p}=\vec{p}^{\perp}+p^{\parallel}\hat{n}_{0}$ with $p=\sqrt{|\vec{p}^{\perp}|^{2}+(p^{\parallel})^{2}}$, suggesting a 2D integral over $\vec{p}^{\perp}$ perpendicular to line-of-sight $\hat{n}_{0}$ and a 1D integral along $\hat{n}_{0}$. Physically speaking, the integral over $p^{\parallel}$ carries the information about the parallel modes of HI field, which is absent in kSZ square field, because kSZ effect is already projected to 2D before taking the square.
@@
with $Y^{(i)}$ defined as three independent non-vanishing contributions from Term $i$ given below
@@
The integral along parallel modes $p^{\parallel}$ appears in the cross-correlation through the convolution when calculating the 2D Fourier transformation of projected HI square field (15), which contains the line-of-sight information that is absent in a pure projected field. Here we have applied a cutoff $p^{\parallel}_{\mathrm{min}}=0.01~\mathrm{Mpc}^{-1}$ accounting for the signal loss due to foregrounds, by which the largest line-of-sight scale mode is removed. Although the cross-correlations with the cutoff value above are almost the same comparing with that with a zero cutoff, it should be reminded that a more aggressive $p^{\parallel}_{\mathrm{min}}$ (such as $0.1~\mathrm{Mpc}^{-1}$) would further suppress the large-scale signal and diminish the expected signal-to-noise ratio.
@@
where each term carries different physical meaning. $C_{\ell}^{(6)}$ is proportional to the cross-correlation between the HI field and the velocity field multiplied by the auto-correlation of the electron density perturbation. Similarly, $C_{\ell}^{(7)}$ encodes the correlation between HI and electron, leaving bulk velocity to be auto-correlated. These two terms correspond to the cross-correlation between HI and the two kinds of perturbation field components of kSZ, i.e. velocity and electron distribution, respectively, and both provide positive contribution to the final cross-correlation. In comparison, $C_{\ell}^{(8)}$ contains cross-correlation of three pairs of perturbation fields: electron-velocity, electron-HI, and HI-velocity, with negative contributions to the final cross-correlation.
II.4 RSD corrections
Since HI intensity mapping is naturally observed in the redshift space, the observed HI intensity has Doppler effect due to the peculiar velocity of the sources. Therefore, the leading-order redshift-space HI density contrast is

### v2

@@ Abstract
Neutral hydrogen (HI) 21-cm intensity mapping is an effective method to track the distribution of baryonic matter, and extract astrophysical and cosmological information. The 21-cm intensity field has a nonvanishing cross-correlation with the kinetic Sunyaev-Zel’dovich (kSZ) effect that traces the velocity and density perturbations of free electrons. By using the linear perturbation theory, in this paper we calculate analytically, for the first time, the cross-correlation between the squared kSZ field and the projection of the squared HI intensity mapping field with the flat-sky approximation. This statistic remains nonvanishing even after the long-wavelength line-of-sight modes ($k_{\parallel}$) are removed due to foreground contamination. We further forecast for the prospects of detection with the SKA-MID 21-cm intensity mapping experiments (redshifts in range of $0.3<z<1$), and the kSZ maps measured by the Atacama Cosmology Telescope (ACT) and Simons Observatory (SO). The predicted cumulative signal-to-noise ratio is $1.92$ for SKA-ACT and $3.99$ for SKA-SO. These results show a possible on-the-edge detection on the cross-correlation signal at low redshifts, which in turn could serve as a validation step toward using it for the Epoch of Reionization studies.
@@ I Introduction
where $\tau$ is the optical depth, and $\sigma_{\mathrm{T}}$ is the Thomson cross-section. $\bar{n}_{\rm e}(z)$ is the mean electron density at redshift $z$, which is equal to
@@
where $\bar{n}_{\mathrm{b},0}$ is the present day’s mean number density of baryons, $X_{\mathrm{e}}(z)$ is the mean ionized fraction at redshift $z$, $\rho_{\rm cr}=1.879h^{2}\times 10^{-29}\,{\rm g}\,{\rm cm}^{-3}$ is the critical density of the Universe, $\Omega_{\rm b}$ is the fractional baryon density, and $\mu_{\rm e}\simeq 1.14$ is the mean mass per electron Ma and Zhao (2014).
@@
In recent years, researchers have considered potential cross-correlations between the patchy reionization-era kSZ signal and other tracers of the Epoch of Reionization (EoR) Hotinli and Johnson (2022); Bégin et al. (2022); Georgiev et al. (2024); Zhou et al. (2025). Among all the large surveys, neutral hydrogen (HI) 21-cm intensity mapping (IM) survey becomes one of the most powerful large-scale structure surveys that can cover wide redshift ranges and large sky areas with efficient scanning strategies Lidz et al. (2011); Battye et al. (2013); Jiang et al. (2023). In recent years, Green Bank Telescope Chang et al. (2010); Masui et al. (2013), Parkes Telescope Pen et al. (2009); Anderson et al. (2018); Tramonte et al. (2019); Tramonte and Ma (2020), and MeerKAT telescope Cunnington et al. (2022) have measured the HI IM to a higher and higher precision, and it is almost certain that the future Square Kilometre Array (SKA) will provide a large survey sample of the HI intensity maps which will constrain cosmological parameters with high precision Santos et al. (2017); Bacon et al. (2020).
@@
In this paper, we consider the cross-correlation between the squared kSZ field and the squared HI intensity field. In each case, it is necessary to square the fields before cross-correlating them. For the kSZ effect, because of the statistical isotropy, clusters with positive and negative peculiar velocities have equal weights. Therefore, the direct two-point correlation function $\langle\delta T_{\mathrm{kSZ}}(\vec{k}_{1})\delta T_{\mathrm{HI}}(\vec{k}_{2})\rangle$ simply vanishes given enough statistics, which can be avoided by taking the square of the kSZ field Dore et al. (2004). In practice, a Wiener filter will be applied before squaring the field to suppress the primary CMB, lensing effects, residual foreground contamination, and detector noise. In the case of the 21-cm field, without squaring, the relevant signal would involve the bispectrum $\langle\delta T_{\mathrm{kSZ}}(\vec{k}_{1})\delta T_{\mathrm{kSZ}}(\vec{k}_{2})\delta T_{\mathrm{HI}}(\vec{k}_{3})\rangle$, which is proportional to $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3})$. Since the kSZ effect is a projected signal, this statistic only receives contributions with $k^{\parallel}_{1}$ and $k^{\parallel}_{2}$ nearly equal to zero, while the Dirac Delta function enforces $k^{\parallel}_{3}=0$, as required by the translational invariance. This means that only the low $k^{\parallel}$ modes in the HI field can contribute to the estimator. However, because the large-scale radial mode of the HI IM is heavily contaminated by the spectrally smooth foregrounds, the HI IM data after foreground removal procedure (either through PCA Davis et al. (1985), FASTICA Wolz et al. (2014), or GNILC Dai and Ma (2025) methods) tend to lose all low-$k^{\parallel}$ modes. Therefore, we need to square the HI IM cube so that we can fully utilize the high-$k^{\parallel}$ modes that are not removed by foreground removal procedure. In Appendix A, we prove that, by taking the square of the HI IM field before projecting it along the line-of-sight we take advantage of all high-$k^{\parallel}$ modes that contribute to the cross-correlation signal. Therefore, our goal is to model cross-correlations between the kSZ-squared field and the projection of the 21-cm-squared field. This is a two-dimensional (2D) cross-correlation which can be performed at the map level.
@@
On large scales, the kSZ effect carries imprints of ionized baryons and follows the distribution of dark matter, which is also the case for the HI field. Therefore, we expect to see a positive correlation function between the two on large scales. On small, galactic scales neutral hydrogen resides predominantly in dense, neutral regions which remain self-shielded against photoionizing radiation Rahmati et al. (2013); Liu et al. (2022). In contrast, free electrons occupy the ionized phase of the circumgalactic medium (CGM), leading to an anticorrelation between the two tracers, which could be seen if the maps have sufficient resolution. However, this small-scale anticorrelation signal may be challenging to see because of the coarse resolution of foreseeable 21-cm measurements. Therefore, we focus our effort in predicting the signal on large scales with $\ell\lesssim 900$, and forecast its detectability.
@@
The paper is organized as follows. In Sec. II, we provide a detailed derivation of the analytical result consisting of the projection method II.1, general formulation of cross-correlations II.2, and the triple-spectrum evaluation II.3. In Sec. III we show our numerical results of the cross-correlation III.1 and the signal-to-noise ratio forecasts III.2 by adopting instrumental parameters of SKA-MID and ACT/SO kSZ measurement. Section IV is devoted to conclusions and discussions.
@@ II Cross-correlation of the projection fields
In this section, we derive an analytical equation for the cross-correlation between the kSZ square field and the HI intensity mapping square field. We first go through the 2D projection procedure of a three-dimensional (3D) field under the flat-sky approximation, and then provide a detailed calculation.
@@ II.1 Two dimension projection field
where $\chi$ is the comoving distance to the observer, $W_{X}$ is the integration kernel for the $X$ field satisfying the normalization condition $\int\mathrm{d}\chi W_{X}(\chi)=1$, and $B_{X}$ is the telescope beam function in the real space which is convolved with the projected field Dore et al. (2004).
@@
Figure 1: An intuitive picture and notation explanation for the flat-sky approximation, with $\chi$ as the comoving distance and $\hat{n}_{0}$ as the line-of-sight. A harmonic decomposition under the flat-sky limit can be approximated to be a 2D Fourier transform over $\vec{\theta}$.
@@
where $B_{\mathrm{HI}}$ denotes the beam factor for HI intensity mapping. In real 21-cm experiments, low-$k^{\parallel}$ modes are often lost due to the foreground removal (foregrounds are smooth in the $k^{\parallel}$ direction), so effectively a filter must be applied to the theory to mimic this observational consequence. In the stage of numerical evaluation, we will apply a low-$k^{\parallel}$ mode cutoff $k^{\parallel}_{\mathrm{min}}=0.01~\mathrm{Mpc}^{-1}$ as the effective filter. $\bar{T}_{\mathrm{b}}(\chi)$ denotes the average brightness temperature at comoving distance $\chi$, which can be evaluated as Harper et al. (2018); Battye et al. (2013)
@@
II.2 Cross-correlating kSZ square field with projection of the HI square field
We define the cross-correlation power spectrum of the kSZ square field $\delta T_{\mathrm{kSZ}^{2}}$ and the projection of the HI intensity mapping square field $\delta T_{\mathrm{HI}^{2}}$ as
@@ II.2 Cross-correlating kSZ square field with projection of the HI square field
The four-point correlation term $\mathcal{T}\equiv\langle q_{z}q_{z}\delta\delta\rangle$ should be proportional to a Dirac delta function by momentum conservation/translation invariance, together with a triple spectrum $\tilde{\mathcal{T}}$ factor
@@
For small angular scales $k^{\parallel}\chi\ll\ell$, one can apply the Limber approximation, allowing us to neglect any $k^{\parallel}$ dependence in the triple spectrum as long as the parallel mode is not enrolled in the convolution; therefore the integral over $k^{\parallel}_{1}$, $k^{\parallel}_{2}$ and $k^{\parallel}_{3}$ can be performed to obtain the Dirac delta functions of comoving distances
@@
which further simplifies the Dirac delta function of perpendicular modes as $\delta_{\mathrm{D}}^{(2)}\left(\vec{\ell}+\vec{\ell}^{\prime}\right)$. Integrating over $\vec{\ell}^{\prime}$ leads to the $C(\vec{\ell})$ expression as follows
@@ II.3 Triple-spectrum calculation
Here we calculate the four-point correlation term appearing in triple-spectrum $\tilde{\mathcal{T}}$, or equivalently the four-point correlator $\mathcal{T}$. An important feature of the kSZ effect Kaiser (1984); Jaffe and Kamionkowski (1998); Ma and Fry (2002) is that most contributions to the signal in integral (1) comes from modes perpendicular to the line-of-sight with $\hat{n}_{0}\cdot\vec{k}=0$ because of (near) cancellations along the line-of-sight direction. Any 3D vector field can be decomposed into a curl-free component ($E$ field) and divergence-free component ($B$ field), $\vec{q}=\vec{q}_{E}+\vec{q}_{B}$, satisfying $\nabla\times\vec{q}_{E}=0$ and $\nabla\cdot\vec{q}_{B}=0$, or in Fourier space $\vec{q}_{E}\propto\hat{k}$ and $\vec{k}\cdot\vec{q}_{B}=0$. Then from the definition given in Eq. (10), the line-of-sight projection can be evaluated as
@@
where have used the fact that the dominant contribution of the kSZ effect comes from the $B$ field.
@@
where $\mu_{1}=\hat{n}_{0}\cdot\hat{p}_{1}$ and $\mu_{2}=\hat{n}_{0}\cdot\hat{p}_{2}$. Further assuming Gaussianity of the perturbation fields allows us to expand the six-point correlation with the Wick Theorem, to a summation of 15 terms in total (see, e.g. Table 1). Before we move into the details of calculation, by observing the underlying symmetry of the arguments in $\mathcal{T}$, some simplifications and reduction can be carried out. First, permutation of $\vec{k}_{3}$ and $\vec{k}_{4}$ does not change the integral, and therefore some of the terms give exactly the same contribution, at the integrand level. For the same reason, permutation of the vector pair $(\vec{k}_{1},\vec{p}_{1})$ and $(\vec{k_{2}},\vec{p}_{2})$ will also result in the same integral because of the exchangeability between $\vec{k}_{1}$ and $\vec{k}_{2}$. As a consequence, all independent terms are shown in Table 1 explicitly, where the power spectrum of any two-point correlation is given by
@@
By applying the arguments replacement Eq.(29), the Dirac delta function $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{3}+\vec{k}_{4})\to\delta_{\mathrm{D}}(k_{3}^{\parallel})\delta_{\mathrm{D}}^{(2)}(\vec{\ell}^{\prime}/\chi_{3})$ indicates a zero contribution to $C_{\ell}$ because it gives out a factor proportional to $\delta_{\mathrm{D}}^{(2)}(\vec{\ell}^{\prime})$. The cross-correlation Eq. (18) has put a constraint $\vec{\ell}+\vec{\ell}^{\prime}=0$ on $\ell$, suggesting that Term $1$ will only contribute to the monopole $\ell=0$, which is not considered. The same factor $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{3}+\vec{k}_{4})$ appears in Terms $2$ and $3$, and thus there is no contribution from these two terms either. Next is Term $4$:
@@
Mathematically, the resultant zero value in the above evaluation is because Term $2$ is proportional to $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{1})$, which does not pick up a specific value of $\vec{p}_{1}$ in the integral, but rather sums over all the directions, leading to a factor proportional to zero. Similarly, Term $5$ is also zero because it is proportional to $\delta_{\mathrm{D}}^{(3)}(\vec{k}_{2})$.
Therefore, the first nontrivial contribution is from Term 66
@@
where the minus sign in the front results from two copies of velocity fields with imaginary parts, as shown in Eq. (11). Note that $k^{\parallel}_{3}$ does not stand for the $k^{\parallel}_{3}$ in Eq. (20) but the $z$-component of $\vec{k}_{3}$. Applying the replacement (29), the triple-spectrum from Term $6$ reads
@@
One can perform a decomposition of convolution momentum $\vec{p}=\vec{p}^{\perp}+p^{\parallel}\hat{n}_{0}$ with $p=\sqrt{|\vec{p}^{\perp}|^{2}+(p^{\parallel})^{2}}$, suggesting a 2D integral over $\vec{p}^{\perp}$ perpendicular to line-of-sight $\hat{n}_{0}$ and a 1D integral along $\hat{n}_{0}$. Physically speaking, the integral over $p^{\parallel}$ carries the information about the parallel modes of the HI field, which is absent in the kSZ square field, because the kSZ effect is already projected to 2D before taking the square.
@@
with $Y^{(i)}$ defined as three independent nonvanishing contributions from Term $i$ given below
@@
The integral along parallel modes $p^{\parallel}$ appears in the cross-correlation through the convolution when calculating the 2D Fourier transformation of the projected HI square field (15), which contains the line-of-sight information that is absent in a pure projected field. Here we have applied a cutoff $p^{\parallel}_{\mathrm{min}}=0.01~\mathrm{Mpc}^{-1}$ accounting for the signal loss due to foregrounds, by which the largest line-of-sight scale mode is removed. Although the cross-correlations with the cutoff value above are almost the same comparing with that with a zero cutoff, it should be reminded that a more aggressive $p^{\parallel}_{\mathrm{min}}$ (such as $0.1~\mathrm{Mpc}^{-1}$) would further suppress the large-scale signal and diminish the expected signal-to-noise ratio.
@@
where each term carries a different physical meaning. $C_{\ell}^{(6)}$ is proportional to the cross-correlation between the HI field and the velocity field multiplied by the autocorrelation of the electron density perturbation. Similarly, $C_{\ell}^{(7)}$ encodes the correlation between HI and the electron, leaving the bulk velocity to be autocorrelated. These two terms correspond to the cross-correlation between HI and the two kinds of perturbation field components of the kSZ, i.e. velocity and electron distribution, respectively, and both provide a positive contribution to the final cross-correlation. In comparison, $C_{\ell}^{(8)}$ contains the cross-correlation of three pairs of perturbation fields: electron-velocity, electron-HI, and HI-velocity, with negative contributions to the final cross-correlation.
II.4 Redshift-space distortion (RSD) corrections
Since HI intensity mapping is naturally observed in the redshift space, the observed HI intensity has the Doppler effect due to the peculiar velocity of the sources. Therefore, the leading-order redshift-space HI density contrast is